# Two tales for a geometric Jensen–Shannon divergence

### v1

@@ Abstract
The geometric Jensen–Shannon divergence (G-JSD) gained popularity in machine learning and information sciences thanks to its closed-form expression between Gaussian distributions. In this work, we introduce an alternative definition of the geometric Jensen–Shannon divergence tailored to positive densities which does not normalize geometric mixtures. This novel divergence is termed the extended G-JSD as it extends to more general positive measures. We give explicitly the gap between the extended G-JSD and G-JSD when considering probability densities, and report both lower and upper bounds in terms of other statistical divergences. We derive corresponding closed-form expressions when considering the case of multivariate Gaussian distributions often met in applications. Finally, we show that these two types of geometric JSDs, the G-JSD and the extended G-JSD, can be interpreted as regularizations of the ordinary JSD by additive terms.
Keywords: Jensen–Shannon divergence; power means; Chernoff information; Bhattacharyya distance; geometric mixtures; exponential families.
@@ 1 Introduction
$D_{\mathrm{KL}}(p_{1},p_{2}){:=}\int p_{1}\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}=E_{p_{1}}\left[\log\frac{p_{1}}{p_{2}}\right].$ (1)
Informally, the KLD quantifies the information lost when $p_{2}$ is used to approximate $p_{1}$ by measuring on average the surprise [^1] Shannon entropy $H(p)=\int p\log\frac{1}{p}\,{\mathrm{d}\mu}$ is the expected surprise $H(p)=E_{p}[-\log p]$ where $\log p(x)$ measures the surprise of the outcome $x$. when outcomes sampled from $p_{1}$ are assumed to emanate from $p_{2}$. Gibbs’ inequality assert that $D_{\mathrm{KL}}(P_{1},P_{2})\geq 0$ if equality if and only if $P_{1}=P_{2}$ $\mu$-almost everywhere. Since $D_{\mathrm{KL}}(p_{1},p_{2})\not=D_{\mathrm{KL}}(p_{2},p_{1})$, various symmetrization schemes of the KLD have been proposed in the literature [6] (e.g., Jeffreys divergence [6, 14], resistor average divergence [15] (harmonic symmetrization), Chernoff information [6], etc.)
@@
$D_{\mathrm{JS}}(p_{1},p_{2}):=\frac{1}{2}\,\left(D_{\mathrm{KL}}(p_{1},a)+D_{\mathrm{KL}}(p_{2},a)\right),$ (2)
where $a=\frac{1}{2}p_{1}+\frac{1}{2}p_{2}$ denotes the statistical mixture of $p_{1}$ and $p_{2}$. The JSD is guaranteed to be upper bounded by $\log 2$ even when the support of $p_{1}$ and $p_{2}$ differ, making it attractive in applications. Furthermore, its square root $\sqrt{D_{\mathrm{JS}}}$ yields a metric distance [9, 31]. The JSD can be extended to a set of densities to measure the diversity of the set as an information radius [35]. In information theory, the JSD can also be interpreted as an information gain [9] since it can be equivalently written as
$D_{\mathrm{JS}}(p_{1},p_{2})=H\left(\frac{1}{2}p_{1}+\frac{1}{2}p_{2}\right)-\frac{H(p_{1})+H(p_{2})}{2},$
where $H(p)=-\int p\log p,{\mathrm{d}\mu}$ is Shannon entropy. [^2] Shannon entropy for discrete measures and differential entropy for continuous measures. The JSD was also defined in the setting of quantum information [5] where it was also proven that its square root yields a metric distance [38].
In [25], the following generalization of the KLD Jensen–Shannon symmetrization scheme was studied for arbitrary statistical dissimilarity $D(\cdot,\cdot)$:
@@
$\mathrm{JS}_{D}^{M_{\alpha},\beta}(p_{1},p_{2}){:=}\beta\,D\left(p_{1},{(p_{1}p_{2})}_{\alpha}^{M}\right)+(1-\beta)\,D\left(p_{2},{(p_{1}p_{2})}_{\alpha}^{M}\right),\quad(\alpha,\beta)\in(0,1)^{2},$ (3)
where ${(p_{1}p_{2})}_{\alpha}^{M}$ is the statistical normalized $M$-mixture of $p_{1}$ and $p_{2}$:
${(p_{1}p_{2})}_{\alpha}^{M}(x){:=}\frac{M_{\alpha}(p_{1}(x),p_{2}(x))}{\int M_{\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}(x)}.$ (4)
@@
(A more general definition is given in [25] by using another arbitrary weighted mean $N_{\beta}$ to average the two KLDs in Eq. 3).
To avoid burden of notations, we write for short $\mathrm{JS}_{D}^{M}(p_{1},p_{2})$ when $\alpha=\beta=\frac{1}{2}$, and further let $D_{\mathrm{JS}}^{M}=\mathrm{JS}_{\mathrm{KL}}^{M}=\mathrm{JS}_{{\mathrm{KL}}}^{M_{\frac{1}{2}},\frac{1}{2}}$.
The ordinary Jensen–Shannon divergence is recovered when $D=D_{\mathrm{KL}}$, $M_{\alpha}(a,b)=\alpha a+(1-\alpha)b$ (weighted arithmetic mean $A_{\alpha}$), and $\beta=\alpha=\frac{1}{2}$. Lower and upper bounds for the skewed $\alpha$-Jensen–Shannon divergence $\mathrm{JS}_{\mathrm{KL}}^{A_{1-\alpha},\frac{1}{2}}$ were reported in [40]. The abstract mixture normalizer of ${(p_{1}p_{2})}_{\alpha}^{M}$ shall be denoted by $Z_{\alpha}^{M}(p_{1},p_{2}){:=}\int M_{\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}(x)$ so that the normalized $M$-mixture is ${(p_{1}p_{2})}_{\alpha}^{M}(x)=\frac{M_{\alpha}(p_{1}(x),p_{2}(x))}{Z_{\alpha}^{M}(p_{1},p_{2})}$.
@@
The generic Jensen–Shannon symmetrization of Definition 1 allows to re-interpret some well-known statistical dissimilarities: For example, the Chernoff information [6, 26] defined by
$D_{C}(p_{1},p_{2}){:=}\max_{\alpha\in(0,1)}D_{B,\alpha}(p_{1},p_{2}),$ (5)
with $D_{B,\alpha}(p_{1},p_{2}){:=}-\log\int p_{1}^{\alpha}p_{2}^{1-\alpha}\,{\mathrm{d}\mu}$ the $\alpha$-skewed Bhattacharrya divergence is a $(\frac{1}{2},\alpha^{*})$ G-$D_{\mathrm{KL}}^{*}$ Jensen–Shannon divergence for the geometric mean $G_{\alpha}(a,b)=a^{\alpha}b^{1-\alpha}$ since for the provably unique optimal skewing factor $\alpha^{*}$ in Eq. 5, we have [26]:
$D_{C}(p_{1},p_{2})=D_{\mathrm{KL}}^{*}(p_{1},(p_{1}p_{2})_{\alpha^{*}}^{G})=D_{\mathrm{KL}}^{*}(p_{2},(p_{1}p_{2})_{\alpha^{*}}^{G}),$
where $D_{\mathrm{KL}}$$^{*}$ is the reverse KLD: $D_{\mathrm{KL}}^{*}(p_{1},p_{2}){:=}D_{\mathrm{KL}}(p_{2},p_{1})$. Hence, it follows that $D_{C}(p_{1},p_{2})=\mathrm{JS}_{{D_{\mathrm{KL}}^{*}}}^{G_{\alpha^{*}},\frac{1}{2}}(p_{1},p_{2})$.
@@
Proposition 1 (G-JSD as Bhattacharyya distance)
We have
$\mathrm{JS}_{{\mathrm{KL}}*}^{G}(p_{1},p_{2})=-\log\int\sqrt{p_{1}p_{2}}\,{\mathrm{d}\mu}=:D_{B}(p_{1},p_{2}),$
where $D_{B}$ denotes the Bhattacharyya distance [4], and
$\mathrm{JS}_{{\mathrm{KL}}*}^{{G_{\alpha}},\alpha}(p_{1},p_{2})=-\log\int p_{1}^{\alpha}p_{2}^{1-\alpha}\,{\mathrm{d}\mu}=:D_{B,\alpha}(p_{1},p_{2}),$
whee $D_{B,\alpha}$ denotes the $\alpha$-skewed Bhattacharyya distance.
@@
Let $m=(p_{1}p_{2})^{G}=\frac{\sqrt{p_{1}p_{2}}}{Z(p_{1},p_{2})}$ be the weighted geometric mixture with normalizer $Z(p_{1},p_{2})=\int\sqrt{p_{1}p_{2}}\,{\mathrm{d}\mu}$. By definition of the JS-symmetrization of the reverse KLD, we have
$\mathrm{JS}_{{\mathrm{KL}}*}^{G}(p_{1},p_{2})$ ${:=}$ $\frac{1}{2}\left(D_{{\mathrm{KL}}^{*}}(p_{1},(p_{1}p_{2})^{G})+D_{{\mathrm{KL}}^{*}}(p_{2},(p_{1}p_{2})^{G})\right),$
$=$ $\frac{1}{2}\left(D_{{\mathrm{KL}}}((p_{1}p_{2})^{G},p_{1},)+D_{{\mathrm{KL}}}((p_{1}p_{2})^{G},p_{2},)\right),$
$=$ $\frac{1}{2}\left(m\,\log\frac{\sqrt{p_{1}p_{2}}}{p_{1}Z(p_{1},p_{2})}+m\,\log\frac{\sqrt{p_{1}p_{2}}}{p_{2}Z(p_{1},p_{2})}\right),$
$=$ $\frac{1}{2}\left(\int\frac{1}{2}m\log\frac{p_{2}}{p_{1}}\,\frac{p_{1}}{p_{2}}{\mathrm{d}\mu}-2\log Z(p_{1},p_{2})\int m\,{\mathrm{d}\mu}\right),$
$=$ $-\log Z(p_{1},p_{2}).$
The proof carries on similarly for the $\alpha$-skewed JS-symmetrization of the reverse KLD: We now let $m_{\alpha}=(p_{1}p_{2})^{G}_{\alpha}=\frac{p_{1}^{\alpha}p_{2}^{1-\alpha}}{Z_{\alpha}(p_{1},p_{2})}$ be the $\alpha$-weighted geometric mixture with normalizer $Z_{\alpha}(p_{1},p_{2})=\int p_{1}^{\alpha}p_{2}^{1-\alpha}\,{\mathrm{d}\mu}$, written as $Z_{\alpha}$ for short below:
$\mathrm{JS}_{{\mathrm{KL}}*}^{G_{\alpha},\alpha}(p_{1},p_{2})$ ${:=}$ $\alpha\,D_{{\mathrm{KL}}^{*}}(p_{1},(p_{1}p_{2})^{G}_{\alpha})+(1-\alpha)\,D_{{\mathrm{KL}}^{*}}(p_{2},(p_{1}p_{2})^{G}_{\alpha}),$
$=$ $\alpha\,D_{{\mathrm{KL}}}(m_{\alpha},p_{1})+(1-\alpha)\,D_{{\mathrm{KL}}}(m_{\alpha},p_{2}),$
$=$ $\int\left(\alpha m_{\alpha}\log\frac{p_{1}^{\alpha}p_{2}^{1-\alpha}}{Z_{\alpha}p_{1}}+(1-\alpha)m_{\alpha}\log\frac{p_{1}^{\alpha}p_{2}^{1-\alpha}}{Z_{\alpha}p_{2}}\right){\mathrm{d}\mu},$
$=$ $-(\alpha+1-\alpha)\log Z_{\alpha}\int m_{\alpha}\,{\mathrm{d}\mu}+\int m_{\alpha}\log\left(\frac{p_{2}}{p_{1}}\right)^{\alpha(1-\alpha)}\,\left(\frac{p_{1}}{p_{2}}\right)^{\alpha(1-\alpha)}{\mathrm{d}\mu},$
$=$ $-\log Z_{\alpha}(p_{1},p_{2}):=D_{B,\alpha}(p_{1},p_{2}).$

### v3

# Frank Nielsen

@@ Abstract
The geometric Jensen–Shannon divergence (G-JSD) gained popularity in machine learning and information sciences thanks to its closed-form expression between Gaussian distributions. In this work, we introduce an alternative definition of the geometric Jensen–Shannon divergence tailored to positive densities which does not normalize geometric mixtures. This novel divergence is termed the extended G-JSD as it applies to the more general case of positive measures. We report explicitly the gap between the extended G-JSD and the G-JSD when considering probability densities, and show how to express the G-JSD and extended G-JSD using the Jeffreys divergence and the Bhattacharyya distance or Bhattacharyya coefficient. The extended G-JSD is proven to be a $f$-divergence which is a separable divergence satisfying information monotonicity and invariance in information geometry. We derive corresponding closed-form formula for the two types of G-JSDs when considering the case of multivariate Gaussian distributions often met in applications. We consider Monte Carlo stochastic estimations and approximations of the two types of G-JSD using the projective $\gamma$-divergences. Although the square root of the JSD yields a metric distance, we show that this is not anymore the case for the two types of G-JSD. Finally, we explain how these two types of geometric JSDs can be interpreted as regularizations of the ordinary JSD.
Keywords: Jensen–Shannon divergence; quasi-arithmetic means; total variation distance; Bhattacharyya distance; Chernoff information; Jeffreys divergence; Taneja divergence; geometric mixtures; exponential families; projective $\gamma$-divergences; $f$-divergence; separable divergence; information monotonicity.
@@ 1.1 Kullback–Leibler and Jensen–Shannon divergences
${\mathrm{KL}}(p_{1},p_{2}){:=}\int p_{1}\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}=E_{p_{1}}\left[\log\frac{p_{1}}{p_{2}}\right].$ (1)
Informally, the KLD quantifies the information lost when $p_{2}$ is used to approximate $p_{1}$ by measuring on average the surprise when outcomes sampled from $p_{1}$ are assumed to emanate from $p_{2}$: Shannon entropy $H(p)=\int p\log\frac{1}{p}\,{\mathrm{d}\mu}$ is the expected surprise $H(p)=E_{p}[-\log p]$ where $\log p(x)$ measures the surprise of the outcome $x$. Logarithms are taken to base $2$ when information is measured in bits, and to base $e$ when it is measured in nats. Gibbs’ inequality assert that ${\mathrm{KL}}(P_{1},P_{2})\geq 0$ with equality if and only if $P_{1}=P_{2}$ $\mu$-almost everywhere. Since ${\mathrm{KL}}(p_{1},p_{2})\not={\mathrm{KL}}(p_{2},p_{1})$, various symmetrization schemes of the KLD have been proposed in the literature [12] (e.g., Jeffreys divergence [12, 25], resistor average divergence [27] (harmonic KLD symmetrization), Chernoff information [12], etc.)
@@
$\mathrm{JS}(p_{1},p_{2}){:=}\frac{1}{2}\,\left({\mathrm{KL}}(p_{1},a)+{\mathrm{KL}}(p_{2},a)\right),$ (2)
where $a=\frac{1}{2}p_{1}+\frac{1}{2}p_{2}$ denotes the statistical mixture of $p_{1}$ and $p_{2}$. The JSD is guaranteed to be upper bounded by $\log 2$ even when the support of $p_{1}$ and $p_{2}$ differ, making it attractive in applications. Furthermore, its square root $\sqrt{\mathrm{JS}}$ yields a metric distance [17, 45].
The JSD can be extended to a set of densities to measure the diversity of the set as an information radius [51]. In information theory, the JSD can also be interpreted as an information gain [17] since it can be equivalently written as
$\mathrm{JS}(p_{1},p_{2})=H\left(\frac{1}{2}p_{1}+\frac{1}{2}p_{2}\right)-\frac{H(p_{1})+H(p_{2})}{2},$
where $H(p)=-\int p\log p\,{\mathrm{d}\mu}$ is Shannon entropy (Shannon entropy for discrete measures and differential entropy for continuous measures). The JSD was also defined in the setting of quantum information [8] where it was also proven that its square root yields a metric distance [55].
Remark 1
Both the KLD and the JSD belong to the family of $f$-divergences [1, 13] defined for a convex generator $f(u)$ (strictly convex at $1$) by:
$I_{f}(p_{1},p_{2}){:=}\int p_{1}\,f\left(\frac{p_{2}}{p_{1}}\right)\,{\mathrm{d}\mu}.$
Indeed, we have ${\mathrm{KL}}(p_{1},p_{2})=I_{f_{\mathrm{KL}}}(p_{1},p_{2})$ and $\mathrm{JS}(p_{1},p_{2})=I_{f_{\mathrm{JS}}}(p_{1},p_{2})$ for the following generators:
$f_{\mathrm{KL}}(u)$ ${:=}$ $-\log u,$
$f_{\mathrm{JS}}(u)$ ${:=}$ $-(1+u)\log\frac{1+u}{2}+u\log u.$
The family of $f$-divergences are the invariant divergences in information geometry [3, 38, 40]. The $f$-divergences guarantee information monotonicity by coarse graining [3] (also called lumping in information theory [14]). Using Jensen inequality, we get that $I_{f}(p_{1},p_{2})\geq f(1)$.
Remark 2
The metrization of $f$-divergences was studied in [46]. Once a metric distance $D(p_{1},p_{2})$ is given, we may use the following metric transform [49] to obtain another metric which is guaranteed to be bounded by $1$:
$0\leq d(p_{1},p_{2})=\frac{D(p_{1},p_{2})}{1+D(p_{1},p_{2})}\leq 1.$
1.2 Jensen–Shannon symmetrization of dissimilarities with generalized mixtures
In [37], a generalization of the KLD Jensen–Shannon symmetrization scheme [39] was studied for arbitrary statistical dissimilarity $D(\cdot,\cdot)$ by using an arbitrary weighted mean [9] $M_{\alpha}$. A generic weighted mean $M_{\alpha}(a,b)=M_{1-\alpha}(b,a)$ for $a,b\in\mathbb{R}_{>0}$ is a continuous symmetric monotonic map $\alpha\in[0,1]\mapsto M_{\alpha}(a,b)$ such that $M_{0}(a,b)=b$ and $M_{1}(a,b)=1$. For example, the quasi-arithmetic means [9] are defined according to a monotonous continuous function $\phi$ as follows:
$M_{\alpha}^{\phi}(a,b){:=}\phi^{-1}\left(\alpha\phi(a)+(1-\alpha)\phi(b)\right).$
[…]
Let us recall the generalization of the Jensen–Shannon symmetrization scheme of a dissimilarity measure presented in [37]:
@@ 1.2 Jensen–Shannon symmetrization of dissimilarities with generalized mixtures
$D^{\mathrm{JS}}_{M_{\alpha},\beta}(p_{1},p_{2}){:=}\beta\,D\left(p_{1},{(p_{1}p_{2})}_{M_{\alpha}}\right)+(1-\beta)\,D\left(p_{2},{(p_{1}p_{2})}_{M_{\alpha}}\right),\quad(\alpha,\beta)\in(0,1)^{2},$ (3)
where ${(p_{1}p_{2})}_{M_{\alpha}}$ is the statistical normalized weighted $M$-mixture of $p_{1}$ and $p_{2}$:
${(p_{1}p_{2})}_{M_{\alpha}}(x){:=}\frac{M_{\alpha}(p_{1}(x),p_{2}(x))}{\int M_{\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}(x)}.$ (4)
Remark 3
@@
A more general definition is given in [37] by using another arbitrary weighted mean $N_{\beta}$ to average the two dissimilarities in Eq. 3:
$D^{\mathrm{JS}}_{M_{\alpha},N_{\beta}}(p_{1},p_{2}){:=}N_{\beta}\left(D\left(p_{1},{(p_{1}p_{2})}_{M_{\alpha}}\right),D\left(p_{2},{(p_{1}p_{2})}_{M_{\alpha}}\right)\right),\quad(\alpha,\beta)\in(0,1)^{2}.$ (5)
When $N_{\beta}=A_{\alpha}$ the weighted arithmetic mean $A_{\alpha}(a,b)=\alpha a+(1-\alpha)b$, Eq. 5 amounts to Eq. 3.
When $\alpha=\frac{1}{2}$, we write for short ${(p_{1}p_{2})}_{M}$ instead of ${(p_{1}p_{2})}_{M_{\frac{1}{2}}}$ in the reminder.
When $D={\mathrm{KL}}$, $M=N=A_{\frac{1}{2}}$, Eq. 5 yields the Jensen–Shannon divergence of Eq. 2: $\mathrm{JS}(p_{1},p_{2})={\mathrm{KL}}^{\mathrm{JS}}_{A_{\frac{1}{2}},A_{\frac{1}{2}}}(p_{1},p_{2})={\mathrm{KL}}^{\mathrm{JS}}_{A,A}(p_{1},p_{2})$.
Lower and upper bounds for the skewed $\alpha$-Jensen–Shannon divergence were reported in [57].
The abstract mixture normalizer of ${(p_{1}p_{2})}_{M_{\alpha}}$ shall be denoted by
$Z_{M_{\alpha}}(p_{1},p_{2}){:=}\int M_{\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}(x),$
so that the normalized $M$-mixture is written as ${(p_{1}p_{2})}_{M_{\alpha}}(x)=\frac{M_{\alpha}(p_{1}(x),p_{2}(x))}{Z_{M_{\alpha}}(p_{1},p_{2})}$. The normalizer $Z_{M_{\alpha}}(p_{1},p_{2})$ is always finite and thus the weighted $M$-mixtures ${(p_{1}p_{2})}_{M_{\alpha}}$ are well-defined:
Proposition 1
For any generic weighted mean $M_{\alpha}$, we have the normalizer of the weighted $M$-mixture bounded by $2$:
$0\leq Z_{M_{\alpha}}(p_{1},p_{2})\leq 2.$
Proof:
Since $M_{\alpha}$ is a scalar weighted mean, it satisfies the following in-betweenness property:
$\min\{p_{1}(x),p_{2}(x)\}\leq M_{\alpha}(p_{1}(x),p_{2}(x))\leq\max\{p_{1}(x),p_{2}(x)\}.$ (6)
Hence, by using the following two identities for $a\geq 0$ and $b\geq 0$:
$\min\{a,b\}$ $=$ $\frac{a+b}{2}-\frac{1}{2}|a-b|,$
$\max\{a,b\}$ $=$ $\frac{a+b}{2}+\frac{1}{2}|a-b|,$
we get
$\int\min\{p_{1}(x),p_{2}(x)\}\,{\mathrm{d}\mu}(x)\leq$ $\int M_{\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}(x)$ $\leq\int\max\{p_{1}(x),p_{2}(x)\}\,{\mathrm{d}\mu}(x),$
$0\leq 1-\mathrm{TV}(p_{1},p_{2})\leq$ $Z_{M_{\alpha}}(p_{1},p_{2})$ $\leq 1+\mathrm{TV}(p_{1},p_{2})\leq 2,$ (7)
where
$\mathrm{TV}(p_{1},p_{2}){:=}\frac{1}{2}\int|p_{1}-p_{2}|\,{\mathrm{d}\mu},$
is the total variation distance, upper bounded by $1$. When the support of the densities $p_{1}$ and $p_{2}$ intersect (i.e., non-singular probability measures $P_{1}$ and $P_{2}$), we have $Z_{M_{\alpha}}(p_{1},p_{2})>0$ and therefore the weighted $M$-mixtures $(p_{1}p_{2})_{M_{\alpha}}$ are well-defined. $\square$
@@
The generic Jensen–Shannon symmetrization of dissimilarities given in Definition 1 allows us to re-interpret some well-known statistical dissimilarities:
For example, the Chernoff information [12, 41] is defined by
$C(p_{1},p_{2}){:=}\max_{\alpha\in(0,1)}B_{\alpha}(p_{1},p_{2}),$ (8)
where $B_{\alpha}(p_{1},p_{2})$ denotes the $\alpha$-skewed Bhattacharrya distance:
$B_{\alpha}(p_{1},p_{2}){:=}-\log\int p_{1}^{\alpha}\,p_{2}^{1-\alpha}\,{\mathrm{d}\mu}$ (9)
When $\alpha=\frac{1}{2}$, we note $B(p_{1},p_{2})=B_{\frac{1}{2}}(p_{1},p_{2})$ the Bhattacharrya distance. Notice that the Bhattacharrya distance is not a metric distance as it violates the triangle inequality of metrics.
Using the framework of JS-symmetrization of dissimilarities, we can reinterpret the Chernoff information as
$C(p_{1},p_{2})=({{\mathrm{KL}}^{*}})^{\mathrm{JS}}_{G_{\alpha^{*}},A_{\frac{1}{2}}}(p_{1},p_{2}),$
where $\alpha^{*}$ is provably the unique optimal skewing factor in Eq. 8 such that we have [41]:
$C(p_{1},p_{2})$ $=$ ${\mathrm{KL}}^{*}(p_{1},(p_{1}p_{2})_{G_{\alpha^{*}}})={\mathrm{KL}}^{*}(p_{2},(p_{1}p_{2})_{G_{\alpha^{*}}}),$
$=$ $\frac{1}{2}\,\left({\mathrm{KL}}^{*}(p_{1},(p_{1}p_{2})_{G_{\alpha^{*}}})+{\mathrm{KL}}^{*}(p_{2},(p_{1}p_{2})_{G_{\alpha^{*}}})\right),$
where $\mathrm{KL}$$^{*}$ denotes the reverse KLD:
${\mathrm{KL}}^{*}(p_{1},p_{2}){:=}{\mathrm{KL}}(p_{2},p_{1}).$
Note that KLD is sometimes called the forward KLD (e.g.,[26]), and we have ${{\mathrm{KL}}^{*}}^{*}(p_{1},p_{2})={\mathrm{KL}}(p_{1},p_{2})$.
@@
Proposition 2 (Bhattacharyya distance as G-JSD)
The Bhattacharyya distance [7] and the $\alpha$-skewed Bhattacharyya distances can be interpreted as JS-symmetrizations of the reverse KLD with respect to the geometric mean $G$:
$B(p_{1},p_{2})$ ${:=}$ $-\log\int\sqrt{p_{1}\,p_{2}}\,{\mathrm{d}\mu}=({\mathrm{KL}}^{*})^{\mathrm{JS}}_{G}(p_{1},p_{2}),$
$B_{\alpha}(p_{1},p_{2})$ ${:=}$ $-\log\int p_{1}^{\alpha}\,p_{2}^{1-\alpha}\,{\mathrm{d}\mu}=({\mathrm{KL}}^{*})^{\mathrm{JS}}_{G_{\alpha}}(p_{1},p_{2}).$
@@
Let $m=(p_{1}p_{2})_{G}=\frac{\sqrt{p_{1}p_{2}}}{Z(p_{1},p_{2})}$ denote the weighted geometric mixture with normalizer $Z_{G}(p_{1},p_{2})=\int\sqrt{p_{1}p_{2}}\,{\mathrm{d}\mu}$. By definition of the JS-symmetrization of the reverse KLD, we have
$({{\mathrm{KL}}^{*}})_{G}^{\mathrm{JS}}(p_{1},p_{2})$ ${:=}$ $\frac{1}{2}\left({{\mathrm{KL}}^{*}}(p_{1},(p_{1}p_{2})_{G})+{{\mathrm{KL}}^{*}}(p_{2},(p_{1}p_{2})_{G})\right),$
$=$ $\frac{1}{2}\left({{\mathrm{KL}}}((p_{1}p_{2})_{G},p_{1})+{{\mathrm{KL}}}((p_{1}p_{2})_{G},p_{2})\right),$
$=$ $\frac{1}{2}\left(\int\left(m\,\log\frac{\sqrt{p_{1}p_{2}}}{p_{1}\,Z_{G}(p_{1},p_{2})}+m\,\log\frac{\sqrt{p_{1}p_{2}}}{p_{2}\,Z_{G}(p_{1},p_{2})}\right)\,{\mathrm{d}\mu}\right),$
$=$ $\frac{1}{2}\left(\int\frac{1}{2}m\log\frac{p_{2}}{p_{1}}\,\frac{p_{1}}{p_{2}}{\mathrm{d}\mu}-2\log Z_{G}(p_{1},p_{2})\int m\,{\mathrm{d}\mu}\right),$
$=$ $-\log Z_{G}(p_{1},p_{2})=:B(p_{1},p_{2}).$
The proof carries on similarly for the $\alpha$-skewed JS-symmetrization of the reverse KLD: We now let $m_{\alpha}=(p_{1}p_{2})_{G_{\alpha}}=\frac{p_{1}^{\alpha}p_{2}^{1-\alpha}}{Z_{G_{\alpha}}(p_{1},p_{2})}$ be the $\alpha$-weighted geometric mixture with normalizer $Z_{G_{\alpha}}(p_{1},p_{2})=\int p_{1}^{\alpha}p_{2}^{1-\alpha}\,{\mathrm{d}\mu}$, written as $Z_{G_{\alpha}}$ for short below:
${{\mathrm{KL}}^{*}}_{G_{\alpha},\alpha}^{\mathrm{JS}}(p_{1},p_{2})$ ${:=}$ $\alpha\,{{\mathrm{KL}}^{*}}(p_{1},(p_{1}p_{2})_{G_{\alpha}})+(1-\alpha)\,{{\mathrm{KL}}^{*}}(p_{2},(p_{1}p_{2})_{G_{\alpha}}),$
$=$ $\alpha\,{{\mathrm{KL}}}(m_{\alpha},p_{1})+(1-\alpha)\,{{\mathrm{KL}}}(m_{\alpha},p_{2}),$
$=$ $\int\left(\alpha m_{\alpha}\log\frac{p_{1}^{\alpha}\,p_{2}^{1-\alpha}}{Z_{G_{\alpha}}\,p_{1}}+(1-\alpha)m_{\alpha}\log\frac{p_{1}^{\alpha}p_{2}^{1-\alpha}}{Z_{G_{\alpha}}\,p_{2}}\right){\mathrm{d}\mu},$
$=$ $-(\alpha+1-\alpha)\log Z_{G_{\alpha}}\int m_{\alpha}\,{\mathrm{d}\mu}+\int m_{\alpha}\log\left(\frac{p_{2}}{p_{1}}\right)^{\alpha(1-\alpha)}\,\left(\frac{p_{1}}{p_{2}}\right)^{\alpha(1-\alpha)}{\mathrm{d}\mu},$
$=$ $-\log Z_{G_{\alpha}}(p_{1},p_{2})=:{B_{\alpha}}(p_{1},p_{2}).$